# Many-body quantum geometry in time-dependent quantum systems with emergent quantum field theory instantaneously

### v2

@@ Abstract
We study many-body quantum geometric effects in time-dependent system with emergent quantum integrable field theory instantaneously. We establish a theorem stating that the Berry connection matrix thus all associated geometric quantities of the system can be precisely characterized by excitations up to two particles from the initial quantum integrable system. To illustrate the many-body geometric influence, we analyze an Ising chain subjected to both a small longitudinal field and a slowly rotating transverse field, whose low-energy physics in the scaling limit is instantaneously governed by the quantum $E_{8}$ integrable field theory. Focusing on the quantum geometric potential (QGP), we show the QGP continuously suppresses the instantaneous energy gaps with decreasing longitudinal field, thereby enhancing many-body Landau-Zener tunneling as evidenced by the Loschmidt echo and its associated spectral entropy. The critical threshold for the longitudinal field strength is determined, where the spectral entropy linearly increases with system size and exhibits hyperscaling behavior when approaching to the threshold. As the longitudinal field passes the threshold and decreases toward zero, the QGP continuously leads to vanishing instantaneous energy gaps involving more low-energy excitations, resulting in increasing spectral entropy indicative of many-body Landau-Zener tunneling. Our results unveil telltale quantum geometric signatures in time-dependent many-body systems, elucidating the intricate interplay between quantum geometry and dynamics.
@@
Due to exact solvability, quantum integrable systems may provide critical insights into quantum geometry in many-body systems. For most quantum integrable systems, the eigenstates and corresponding transition matrix elements of physical observables can be calculated exactly using systematic methods, such as the Bethe Ansatz  [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or/and form factor bootstrap [19, 20, 21, 22, 23]. Over the past decade, significant research has focused on quantum integrability in non-equilibrium systems, including integrable quantum quenches and integrable brickwork circuits [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. These studies have provided new insights into non-equilibrium physics from the perspective of integrability, with potential connections to emerging fields such as quantum computing and quantum information. However, the understanding of quantum geometric effects in quantum many-body systems remains underdeveloped [37].
To study quantum geometry and its effects in quantum many-body systems, we focus on time-dependent systems with emergent quantum integrable field theory instantaneously. The $n^{th}$ instantaneous eigenfunction is $|\phi_{n}(t)\rangle$ with instantaneous eigenenergy $e_{n}(t)$ at the moment $t$. After imposing a proper constraint on the adiabatic transformation $\mathscr{U}_{a}(t):|{\phi}_{n}(0)\rangle\to|{\phi}_{n}(t)\rangle$, we prove that the Berry connection matrix (BCM) entries $\gamma_{nm}=i\langle\phi_{n}(t)|\dot{\phi}_{m}(t)\rangle$ only associate excitations up to two particle processes.
@@
Figure 1: An illustration of an eigenstate in an IFT containing $n$ stable quasiparticles on a cylinder of a circumference $L$, with the adiabatic transformation $\mathscr{U}_{a}(t)$.
@@
The BCM in instantaneously quantum-integrable time-dependent systems.— It is generally difficult to calculate the BCM in interacting quantum many-body systems. We consider time-dependent quantum system with emergent 1+1D quantum field theory instantaneously, where the initial state is an eigenstate in an IFT containing $n$ stable quasiparticles, and the instantaneous eigenstate is connected to the initial state with an adiabatic transformation, as $|\phi_{n}(t)\rangle=\mathscr{U}_{a}(t)|\phi_{n}(0)\rangle_{\text{init},L}$. Without loss of generality the system is placed on a cylinder with circumference $L$. We require the effective adiabatic Hamiltonian [47] $\mathcal{H}_{adia}=i\mathscr{U}_{a}^{-1}(t)\dot{\mathscr{U}_{a}}(t)$ to satisfy
[…]
@@
in the continuum limit. Here $\mathcal{O}_{i}(x)$ are the local field densities. The initial state is given by $|\phi_{n}(0)\rangle_{\text{init},L}=|A_{1}(\vartheta_{1};I_{1})...A_{N}(\vartheta_{N};I_{N})\rangle_{L}$. Here, $A_{j},j=1,...,N$ label different types of particles with corresponding mass $m_{j}$, discrete quantum number $I_{j}$, and rapidity $\vartheta_{j}$. In addition, the energy and momentum of this state are given by $E=\sum_{j}m_{j}\cosh\vartheta_{j}$ and $P=\sum_{j}m_{j}\sinh\vartheta_{j}$. For any single particle $A_{j}(\vartheta_{j};I_{j})$, the discrete momentum and rapidity are connected with each other by the Bethe-Yang quantization condition derived from the spatial periodic boundary condition [48, 49, 50, 35, 42]. An illustration is shown in Fig. 1. The BCM can then be explicitly given with the help of quantum integrability. Considering two such states containing $n$ and $m$ quasiparticles respectively, for the off-diagonal matrix element $m\neq n$, the BCM is given by
@@
$\frac{\gamma_{nm}}{L}\approx\sum_{i}f_{i}(t)\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi,\vartheta_{1},...,\vartheta_{M})}{\sqrt{\rho_{N}(\vartheta^{\prime}_{1},...,\vartheta^{\prime}_{N})}\sqrt{\rho_{M}(\vartheta_{1},...,\vartheta_{M})}},$ (2)
@@
with Lüscher’s remainder $o(e^{-\mu L})$ [51, 48, 49, 50]. Here, the function $\mathcal{F}_{nm}^{\mathcal{O}_{i}}$ in terms of rapidities is known as the form factor for certain IFT, and the $\rho$ function comes from the Jacobian of the particle density of states determined by the Bethe-Yang condition. For the diagonal matrix element $m=n$, the result is given by
$\frac{\gamma_{nn}}{L}$ $\approx\sum_{i}f_{i}(t)\bigg{[}{}_{L}\langle\mathcal{O}_{i}\rangle_{L}+$ (3)
@@
$\sum_{A\in\{A_{1},...,A_{N}\}}\frac{\mathcal{F}_{NN}^{\mathcal{O}_{i}}(A)\rho(\{A_{1},...,A_{N}\}\setminus A)_{L}}{\rho(\{A_{1},...,A_{N}\})_{L}}\bigg{]},$
@@
with the Lüscher’s remainder as well. The full information of the BCM implies that for such quantum many-body systems, we can determine all the geometric quantities associated with the BCM. Notably, the Jacobians in the denominator provide suppression behavior of order $L^{n}$ with $n$ being the number of particles involved. Consequently, after simple power counting, we can immediately observe that the excitations involving more than two particles do not contribute to the BCM in the thermodynamic limit ($L\rightarrow\infty$). Thus we prove the following theorem.
@@
Theorem: Consider time-dependent many-body quantum systems with emergent quantum integrable field theory instantaneously, if the effective adiabatic Hamiltonian follows Eq. (1), then in the thermodynamic limit the corresponding BCM entries can be analytically obtained as excitations up to two particles.
In the following, the theorem is applied to a time-dependent quantum Ising model, where the QGP constructed from the BCM entries is shown to play a crucial role in the MBLZT in the system.
@@
[…]
@@
The spectral entropy of the Loschmidt echo.— In order to investigate the role of the EEG in the MBLZT process of a time-dependent quantum system, a commonly adopted approach is to examine the fidelity between the time-evolving wavefunction and the instantaneous eigenstates, denoted as ${}_{ins}\langle\psi_{n}(t)|\psi(t)\rangle$, which is known as the Loschmidt echo (LE).
@@
[…]
@@
Conclusions.— In this article we lay down a theorem revealing the quantum geometry in time-dependent many-body systems whose low-energy can be organized by quantum integrable field theory instantaneously. The theorem states that the BCM entries for local field densities in the system are governed by at most two-particle excitations. The result provides a geometric foundation for diabatic time evolutions mediated by the vanishing EEG due to the QGP contribution. We then analyze a time-dependent quantum Ising chain, showing that the suppression of the EEG with decreasing longitudinal field enhances the MBLZT. Near the critical ratio $\kappa=1$, the hyperscaling behavior of the LE spectral entropy signals a sharp transition in the MBLZT. For $\kappa<1$, the numerical results show persistent entropy growth mixing with a cascade of peaks, reflecting continuous enhancement of MBLZT. Our results indicate that the LE spectral entropy can serve as a powerful diagnostic tool for diabatic time evolution and its associated MBLZT. Yet, a full physical understanding on the analytical structure of the LE spectral entropy, such as its hyperscaling behavior, linear in $L$ behavior at $\kappa=1$, and additional peaks when $\kappa<1$, requires more analytical and numerical efforts, a challenge that remains open but worth to explore further. Meanwhile it is also worth to generalize the theorem in quantum integrable lattice systems, and extend to the BCM entries for quasi-local field densities.
The time-dependent Hamiltonian in Eq. (4) is experimentally accessible in platforms such as Rydberg atomic arrays, [55, 56, 35] a promising setup for direct experimental verification of our theoretical predictions. Furthermore, our study reveals the important role of quantum geometry in time-dependent quantum many-body systems, providing an important insight in understanding non-equilibrium many-body systems.

### v3

# Berry connection and quantum geometry in time-dependent systems with instantaneous quantum integrable field theory

@@ Abstract
We study many-body quantum geometric effects in time-dependent system with emergent quantum integrable field theory instantaneously. We establish a theorem stating that the Berry connection matrix thus all associated geometric and topological quantities of the system can be precisely characterized by excitations up to two particles from the initial emergent quantum integrable field theory. To illustrate the many-body geometric influence, we analyze an Ising chain subjected to both a small longitudinal field and a slowly rotating transverse field, whose low-energy physics in the scaling limit is instantaneously governed by the quantum $E_{8}$ integrable field theory. Focusing on the quantum geometric potential (QGP), we show the QGP continuously suppresses the instantaneous energy gaps with decreasing longitudinal field, thereby enhancing many-body Landau-Zener tunneling as evidenced by the Loschmidt echo and its associated spectral entropy. The critical threshold for the longitudinal field strength is determined, where the spectral entropy linearly increases with system size and exhibits hyperscaling behavior when approaching to the threshold. As the longitudinal field passes the threshold and decreases toward zero, the QGP continuously leads to vanishing instantaneous energy gaps involving more low-energy excitations, resulting in increasing spectral entropy indicative of many-body Landau-Zener tunneling. Our results unveil telltale quantum geometric signatures in time-dependent many-body systems, elucidating the intricate interplay between quantum geometry, integrability, and dynamics.
@@
Due to exact solvability, quantum integrable systems may provide critical insights into quantum geometry in many-body systems: Their eigenstates can be completely labeled by the corresponding conserved quantum numbers [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], therefore provide a concrete base to analytically study quantum geometry in quantum integrable systems [21, 22, 23, 24, 25, 26]. Over the past decade, significant research has focused on quantum integrability in non-equilibrium systems, including integrable quantum quenches and integrable brickwork circuits [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. These studies have provided new insights into non-equilibrium physics from the perspective of integrability, with potential connections to emerging fields such as quantum computing and quantum information. However, the understanding of quantum geometric effects in quantum many-body (integrable) systems remains underdeveloped [40].
To study quantum geometry and its effects in quantum many-body systems, we focus on time-dependent systems with emergent quantum integrable field theory instantaneously. The $n$-th instantaneous eigenfunction is $|\phi_{n}(t)\rangle$ with instantaneous eigenenergy $e_{n}(t)$ at the moment $t$. After imposing a proper constraint on the adiabatic transformation $\mathscr{U}_{a}(t):|{\phi}_{n}(0)\rangle\to|{\phi}_{n}(t)\rangle$, we prove that the Berry connection matrix (BCM) entries $\gamma_{nm}=i\langle\phi_{n}(t)|\dot{\phi}_{m}(t)\rangle$ only associate excitations up to two particle processes.
@@
Figure 1: An illustration of the setup for the theorem. (a) The initial wavefunction of an eigenstate in an IFT contains NN quasiparticles with periodic boundary conditions, represented as a cylinder with circumference $L$. The blue circles represent the quasiparticles. (b) The time evolution of the initial wavefunction under the unitary transformation $\mathscr{U}_{a}(t)$, with the effective adiabatic Hamiltonian given by Eq. (1). The blue line corresponds to the initial configuration in (a), and the red space-time network illustrates $\mathscr{U}_{a}(t)$.
@@
The BCM in instantaneously quantum-integrable time-dependent systems.— It is generally difficult to calculate the BCM in interacting quantum many-body systems. We consider time-dependent quantum system with emergent 1+1D quantum field theory instantaneously, where the initial state is an eigenstate in an IFT containing $N$ stable quasiparticles [Fig. 1(a)], and the instantaneous eigenstate is connected to the initial state with an adiabatic transformation, as $|\phi_{N}(t)\rangle=\mathscr{U}_{a}(t)|\phi_{N}(0)\rangle_{\text{init},L}$ [Fig. 1(b)]. Without loss of generality the system is placed on a cylinder with circumference $L$. We require the effective adiabatic Hamiltonian [48] $\mathcal{H}_{adia}=i\mathscr{U}_{a}^{-1}(t)\dot{\mathscr{U}_{a}}(t)$ to satisfy
$\mathcal{H}_{adia}=\sum_{i=1}^{N_{\mathcal{O}}}\int_{0}^{L}f_{i}(t)\mathcal{O}_{i}(x)dx$ (1)
@@
in the continuum limit. Here $\mathcal{O}_{i}(x)$ are the local field densities. The initial state is given by $|\phi_{N}(0)\rangle_{\text{init},L}=|A_{1}(\vartheta_{1};I_{1})...A_{N}(\vartheta_{N};I_{N})\rangle_{L}$. Here, $A_{j},j=1,...,N$ label different types of particles with corresponding mass $m_{j}$, discrete quantum number $I_{j}$, and rapidity $\vartheta_{j}$. In addition, the energy and momentum of this state are given by $E=\sum_{j}m_{j}\cosh\vartheta_{j}$ and $P=\sum_{j}m_{j}\sinh\vartheta_{j}$. For any single particle $A_{j}(\vartheta_{j};I_{j})$, the discrete momentum and rapidity are connected with each other by the Bethe-Yang quantization condition derived from the spatial periodic boundary condition [49, 50, 51, 38, 43]. The BCM can then be explicitly given with the help of quantum integrability. Considering two such states containing $N$ and $M$ quasiparticles respectively, for the off-diagonal matrix element where $M\neq N$, or $M=N$ but the two sets of rapidities are not exactly same, the BCM is given by
@@
$\frac{\gamma_{NM}}{L}\approx\sum_{i}f_{i}(t)\frac{\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\{\vartheta^{\prime}+i\pi\}_{N},\{\vartheta\}_{M})}{\sqrt{\rho_{L,\{\mathscr{A^{\prime}}\}_{N}}}\sqrt{\rho_{L,\{\mathscr{A}\}_{M}}}},$ (2)
@@
with Lüscher’s remainder $o(e^{-\mu L})$ [52, 49, 50, 51]. Here, $\{\vartheta^{\prime}+i\pi\}_{N}=\{\vartheta^{\prime}_{1}+i\pi,...,\vartheta^{\prime}_{N}+i\pi$} and $\{\vartheta\}_{M}=\{\vartheta_{1},...,\vartheta_{M}\}$ are the sets of rapidities of the two particle sets $\{\mathscr{A^{\prime}}\}_{N}=\{A^{\prime}_{1},...,A^{\prime}_{N}\}$ and $\{\mathscr{A}\}_{M}=\{A_{1},...,A_{M}\}$. The function $\mathcal{F}_{NM}^{\mathcal{O}_{i}}(\{\vartheta^{\prime}+i\pi\}_{N},\{\vartheta\}_{M})$ in terms of rapidities is the matrix element of operator $\mathcal{O}_{i}$ with respect to the two multi-particle states, known as the form factor of IFT [43, 21, 20, 22, 23]. The two functions of $\rho_{L,\{\mathscr{A^{\prime}}\}_{N}}$ and $\rho_{L,\{\mathscr{A}\}_{M}}$ are Jacobian from the particle density of states of the two particle sets in the finite size system, where the detailed expressions can be found in [43]. For the diagonal matrix element where both $M=N$, and the two sets of rapidities are exactly the same, the result is given by
$\frac{\gamma_{NN}}{L}$ $\approx\sum_{i}f_{i}(t)\bigg[{}_{L}\langle\mathcal{O}_{i}\rangle_{L}$ (3)
@@
$+\sum_{\forall\mathscr{S}\subseteq\{\mathscr{A}\}_{N}}\frac{\mathcal{F}_{\mathscr{S}\mathscr{S}}^{\mathcal{O}_{i}}(\{\vartheta+i\pi\}_{\mathscr{S}},\{\vartheta\}_{\mathscr{S}})\rho_{L,\{\mathscr{A}\}_{N}\setminus\mathscr{S}}}{\rho_{L,\{\mathscr{A}\}_{N}}}\bigg],$
@@
with the Lüscher’s remainder as well. Here $_{L}\langle\mathcal{O}_{i}\rangle_{L}$ corresponds to the vacuum expectation value of $\mathcal{O}_{i}$, $\mathscr{S}$ is non-empty subset of $\{\mathscr{A}\}_{N}$. $\{\mathscr{A}\}_{N}\setminus\mathscr{S}$ denotes particle set excluding $\mathscr{S}$ from $\{\mathscr{A}\}_{N}$. The full information of the BCM implies that for such quantum many-body systems, we can determine all the geometric quantities associated with the BCM. Notably, the Jacobians in the denominator provide suppression behavior of order $L^{\mathcal{Z}}$ with total number of particle $\mathcal{Z}=M+N$. The algebraic decaying behaviour of the Jacobians in the limit of $L\rightarrow\infty$ [43] consequently leads to the asymptotic behaviour of the BCM elements. Thus we prove the following theorem.
@@
Theorem: Consider time-dependent many-body quantum systems with emergent quantum integrable field theory instantaneously, if the effective adiabatic Hamiltonian follows Eq. (1), then the corresponding BCM entries with total quasi-particle numbers of $\mathcal{Z}$ follow the asymptotic algebraic behaviour $\sim L^{1-\mathcal{Z}/2}$ as the system size $L\rightarrow\infty$.
This theorem suggests that in the thermodynamic limit ($L\rightarrow\infty$) the corresponding BCM entries can be analytically determined as excitations up to two particles. This can also be understood from the quasi-particle nature of IFTs: While integrating the field operator over the entire system introduces a volume dependence, the Jacobian factor arising from the density of particle states ultimately suppresses the asymptotic behaviour of the BCM elements for multi-particle states. In the following, the theorem is applied to a time-dependent quantum Ising model, where the QGP constructed from the BCM entries is shown to play a crucial role in the MBLZT in the system.
@@
In the scaling limit, for the low-energy sector the EEG during the time evolution can be solved analytically. Considering any two instantaneous eigenstates $|\psi_{n}(t)\rangle_{ins}$ and $|\psi_{m}(t)\rangle_{ins}$ in the model Eq. (4), the instantaneous energy gap reads $E_{mn}=\sum_{\beta=1}^{M}m_{\beta}\cosh\theta_{\beta}-\sum_{\alpha=1}^{N}m_{\alpha}\cosh\theta^{\prime}_{\alpha}$, where we have assumed that there are totally $M$ and $N$ quantum $E_{8}$ particles in the two states, respectively. With the mass ratios for the $E_{8}$ particles $\xi_{i}$ in units of $m_{1}$, the masses of the quantum $E_{8}$ particles follows $m_{\beta(\alpha)}=\xi_{\beta(\alpha)}m_{1}=\xi_{\beta(\alpha)}C_{m}h^{8/15}_{z}$, and $\theta^{(^{\prime})}_{\beta(\alpha)}$ is the corresponding rapidity. The BCM entry is given by $\gamma_{km}(t)=-\dot{\varphi}_{m}(t)\delta_{km}-\omega_{0}/2\int dx\langle\psi_{k}|\sigma(x)|\psi_{m}\rangle$. Applying the quantum $E_{8}$ form factor theory, the QGP can be organized in a more compact form as [43]
@@
[…]
@@
We use the TLFFA algorithm to calculate the quantities introduced above. The details of the TLFFA algorithm can be found in [43]. Specifically, in this algorithm, we truncate the free fermion basis to an energy of 10JJ above the vacuum. For different system sizes, we ensure convergent results by increasing the number of states. For instance, for $L=100$ the number truncation of states reaches $16384$ for convergence, and this also reaches our computational limit.
The results are summarized in Fig. 3. When $\kappa\gg 1$, $S_{\text{LE}}$ is small, indicating weak MBLZT. As $\kappa$ decreases, $S_{\text{LE}}$ increases continuously, suggesting an enhancement of MBLZT. When $\kappa=1$, a peak appears and increases linearly in $L$ [Fig. 3 inset (a)], which possibly implies MBLZT avalanche in thermodynamic limit. As $\kappa$ decreases further below $8/15$ upper bound of zeros of the EEG between the ground state and excited states, it is expected that the MBLZT is further strengthened, as manifested in the continuously increasing $S_{\text{LE}}$. Additional peaks appear when $\kappa<1$, signaling further strengthened MBLZT at specific $\kappa$’s. For $\kappa\rightarrow 0^{+}$, it is noticed that the increase of $S_{\text{LE}}$ becomes weaker for $L=100$ comparing to smaller $L$ [Fig. 3 (b)], which possibly is due to the truncated energy cut at 10JJ in our numerical calculation. A higher cut should provide better results, yet it exceeds our computational limit. Interestingly with $\kappa$ approaching to 1, $S_{\text{LE}}$ exhibits a hyperscaling behavior in terms of the scaling parameter $(\kappa-1)^{8/15}JL$ [Fig. 3 inset (b), $J=1$], implying emergence of universal physics near $\kappa=1$.
@@
Conclusions.— In this article we lay down a theorem revealing the quantum geometry in time-dependent quantum many-body systems whose low-energy can be organized by quantum integrable field theory instantaneously. The theorem states that the BCM entries for local field densities in the system are governed by at most two-particle excitations. The result provides a geometric foundation for diabatic time evolutions mediated by the vanishing EEG due to the QGP contribution.
We then analyze a time-dependent quantum Ising chain, showing that the suppression of the EEG with decreasing longitudinal field enhances the MBLZT. Near the critical ratio $\kappa=1$, the hyperscaling behavior of the LE spectral entropy signals a sharp transition in the MBLZT. For $\kappa<1$, the numerical results show persistent entropy growth mixing with a cascade of peaks, reflecting continuous enhancement of MBLZT. Our results indicate that the LE spectral entropy can serve as a powerful diagnostic tool for diabatic time evolution and its associated MBLZT. Yet, a full physical understanding on the analytical structure of the LE spectral entropy, such as its hyperscaling behavior, linear in $L$ behavior at $\kappa=1$, and additional peaks when $\kappa<1$, requires more analytical and numerical efforts, a challenge that remains open but worth to explore further. Meanwhile it is also worth to generalize the theorem in quantum integrable lattice systems, and extend to the BCM entries for quasi-local field densities.
The time-dependent Hamiltonian in Eq. (4) is experimentally accessible in platforms such as Rydberg atomic arrays, [56, 57, 38] a promising setup for direct experimental verification of our theoretical predictions.